INFRARED ABSORPTION OF CH_{3}SONO DETECTED WITH TIME-RESOLVED FOURIER-TRANSFORM SPECTROSCOPY

\underline{YUAN-PERN LEE}, Department of Applied Chemistry and Institute of Molecular Science, National Chiao Tung University, Hsinchu 30010, Taiwan and Institute of Atomic and Molecular Sciences, Academia Sinica, Taipei 10617, Taiwan; JIN-DAH CHEN, Department of Applied Chemistry, National Chiao Tung University, Hsinchu 30010, Taiwan.

A step-scan Fourier-transform spectrometer coupled with a 6.4-m multipass absorption cell was employed to detect timeresolved infrared absorption spectra of reaction intermediates produced upon UV irradiation of a flowing mixture of CH_{3}SSCH_{3} and NO_{2} in CO_{2}. Irradiation of CH_{3}SSCH_{3} at 248 nm produces CH_{3}S radicals that subsequently react with NO_{2}. Under a total pressure of 100 Torr, we observed bands near 1560 cm^{−1}, assignable to mainly the N=O stretching mode of CH_{3}SONO, with a small contribution from CH_{3}SNO_{2}. Calculations with density-functional theory (B3LYP/aug-cc-pVTZ and B3P86/aug-cc-pVTZ) predicted the geometry, vibrational wavenumbers, and rotational parameters of CH_{3}SONO and CH_{3}SNO_{2}. Based on these predicted rotational parameters, the simulated absorption band agrees satisfactorily with experimental results. Under a total pressure of 16 Torr, bands near 1560 and 1260 cm^{−1} are assigned to NO_{2} asymmetric and symmetric stretching modes of CH_{3}SNO_{2}, respectively; the former is overlapped with the N=O stretching mode of CH_{3}SONO. An additional band near 1070 cm^{−1} is assigned to the S=O stretching mode of CH_{3}SO, reported previously as a secondary product in the reaction of CH_{3}S + O_{2}.^{a} Reaction of CH_{3}S + NO_{2} at high pressure clearly yields CH_{3}SONO, rather than CH_{3}SNO_{2}, as a major product.

aL.-K. Chu and Y.-P. Lee, J. Chem. Phys. \underline{133}, 184303 (2010).